If 51% of the population carries at least one copy of the recessive allele. what is the predicted frequency of the population expressing the dominant phenotype?



1
Expert's answer
2022-03-18T06:50:01-0400

According to Hardy-Weinberg equations p+q=1 and p2+2pq+q2=1

Where p=frequency of dorminant allele

q=frequency of recessive allele

p2= frequency of heterozygous genotype

q2= frequency of recessive genotype

Since we do not have the frequency, we will first determine the value of p and q.

P2 would be equal to 100-51%=49%

So p=0.7

Q=1-p. =1-0.7=0.3

The predicted frequency of individuals who express the dorminant phenotype are frequency of homozygous dorminant+ the frequency of heterozygous or p2+2pq

=0.72+(2*0.7*0.3)

=0.49+0.42

=0.91

So the correct answer is 0.91